# Graph-Informed Neural Networks for Sparse Grid-Based Discontinuity Detectors

### v1

@@ 3 Discontinuity Detection and Sparse Grids
The discontinuity detection method proposed herein aims to extend the one proposed in [30] moving from uniform grids to sparse grids. In view of this extension, we seek for a NN architecture leveraging sparse grids; thus, we consider the use of Graph-Informed Neural Networks (GINNs) built with respect to a SGG. We remark that the method proposed herein is designed in such a way that also classic NNs, such as Multi-Layer Perceptrons (MLPs), can be used; however, the numerical experiments show that the method benefits from the use of GINN (see Section 5); this is in line with the results presented in [5].
For a given function $f:\mathbb{R}^{n}\to\mathbb{R}$, the method proposed relies on the following ingredients:
  • the concept of trouble points of a sparse grid (i.e, points of the grid that are near to a discontinuity interface of $f$);
@@
The algorithm defined in this work does not depend on the nature of the discontinuity detector used, and the function $f$ can be rather general; however, the algorithm works better with discontinuous functions characterized by discontinuity interfaces with small codimensions (with respect to $\mathbb{R}^{n}\supseteq\Omega$. In the following, we formally introduce the notions of troubled point and discontinuity detector based on a sparse grid. The definition of troubled points is inspired by the definition of troubled cells given in [30].
@@
Let $\mathcal{S}$ be a sparse grid in $\mathbb{R}^{n}$ and let $G=(\mathcal{S},E)$ be the (possibly weighted) graph based on $\mathcal{S}$. Let $g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a given function such that $\mathbb{B}(\mathcal{S})\subseteq\Omega$ and let $\delta$ be the set of all and only the discontinuity points of $g$. Then, the node $\boldsymbol{x}_{i}$ of $\mathcal{S}$ is said to be a troubled point with respect to $g$ if the following conditions are satisfied. There exist $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}\in E$ such that:
  • $\delta\cap e_{ij}\neq\emptyset\,$;
@@
Let $\mathcal{S}$ be a sparse grid in $\mathbb{R}^{n}$ made of $N$ points and let $\mathcal{S}_{/\simeq}$ denotes the set of all the sparse grids similar to $\mathcal{S}$. Let $\mathcal{F}$ denotes the set of all the functions $g:\Omega\rightarrow\mathbb{R}$, for an arbitrary $\Omega\subseteq\mathbb{R}^{n}$. Then, a discontinuity detector based on $\mathcal{S}$ is a function $\Delta$ that, for each $\mathcal{S}^{\prime}\simeq\mathcal{S}$ and each function $g\in\mathcal{F}$, returns a vector $\boldsymbol{p}=(p_{1},\ldots,p_{N})$, with $0\leq p_{i}\leq 1$ for all $i=1,\ldots,N$, where $p_{i}$ gives an indication of how likely $\boldsymbol{x}_{i}^{\prime}\in\mathcal{S}^{\prime}$ is a troubled point.
According to 3.2, a discontinuity detector provides, for each grid point, a sort of likelihood for being a troubled point. A point $\boldsymbol{x}_{i}^{\prime}$ is labeled as troubled if $p_{i}\geq\tau$, where $\tau\in(0,1]$ is a given threshold.
The previous definition is intentionally qualitative, as several types of discontinuity detectors can be defined. For example, we can define detectors that evaluate $\boldsymbol{p}$ knowing: $i$) the function $g$ and the coordinates of the points in $\mathcal{S}^{\prime}$; $ii$) the evaluations of $g$ at the points of $\mathcal{S}^{\prime}$ and the points' coordinates; $iii$) the evaluations of $g$ at the points of $\mathcal{S}^{\prime}$, only. In this work, we focus on building NN-based detectors of the latter type; i.e., detectors with function $\Delta$ such that $\Delta:(\mathbb{R}\cup\{\infty\})^{N}\rightarrow[0,1]^{N}$ and
@@
However, an exact detector is typically defined through expensive algorithms and in general it is not a detector based on the discontinuous function evaluations only (see (11)). Therefore, given a sparse grid, our idea consists in building a non-expensive but effective discontinuity detector that is only-evaluations dependent, training a NN model for approximating the exact one; given such a kind of NN-based detector, we can use it with a sparse grid-based discontinuity detection algorithm, aiming to reach very good results at a consistently lower computational cost, even in domains of dimension $n>3$.
Figure 3: Example of sparse grid points classification made for a function $g:\mathbb{R}^{2}\to\mathbb{R}$ with an exact discontinuity detector $\Delta^{*}$ (left) and a inexact discontinuity detector $\Delta$ (right). The black dotted line is the function's discontinuity interface. Green circles are true troubled points, magenta circles are false troubled points, magenta dots are false non-troubled points. For the inexact detector, we report also the values $p_{i}$ corresponding to the highlighted sparse grid points (threshold $\tau=0.5$).
@@
Similarly to what done in [30], the introduction of a NN-based detector that is only-evaluations dependent (see (11)) is useful to make the detector universally applicable to any sparse grid $\mathcal{S}^{\prime}\simeq\mathcal{S}$ and any discontinuous function with domain dimension equal to $n$. On the other hand, since these detector will not be exact, we can expect that the smaller are the ratios between the function evaluations and the edge lengths (i.e., the slopes $|g(\boldsymbol{x}_{i}^{\prime})-g(\boldsymbol{x}_{j}^{\prime})|/\|{\boldsymbol{x}_{i}^{\prime}-\boldsymbol{x}_{j}^{\prime}}\|$), the lower is the reliability of the predicted values for $\boldsymbol{p}$.
@@ 3.1 Algorithm for Discontinuity Detection
We adopt a similar approach for our sparse grid-based method. However, thanks to 2.1 and 3.1, we can use the same detector $\Delta$ for both the coarse step and any refinement step. Therefore, we define an iterative approach that stops when a minimum box size for the sparse grid is reached. We point the reader to the fact that we can apply several refinements leveraging the use of sparse grids, which allow to save on the number of function evaluations, thus delaying the curse of dimensionality phenomenon.
@@
    using $\Delta$, detect the troubled points of one sparse grid and mark it as ``checked'';
@@
    for each troubled point $\boldsymbol{x}$, generate a new sparse grid $\mathcal{S}^{\prime}$ similar to $\mathcal{S}$, with box $\mathbb{B}(\mathcal{S}^{\prime})$ centered in $\boldsymbol{x}$ and box's edge length equal to the length of the longest graph's edge intersecting the troubled point. If this latter length is shorter than a chosen value $\Lambda_{\min}\in\mathbb{R}_{+}$, do not create the new sparse grid.
@@
    repeat steps 2 and 3, until all the sparse grids are ``checked''.
@@
1:$g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$, function; $\mathcal{S}$, sparse grid (equispaced, hypercubic box); $\Delta$, discontinuity detector based on $\mathcal{S}$; $L=((\boldsymbol{x}_{1},\lambda_{1}),\ldots,(\boldsymbol{x}_{m},\lambda_{m}))$, sequence of center-edge length pairs characterizing the sparse grid boxes $\mathbb{B}(\mathcal{S}^{(1)}),\ldots,\mathbb{B}(\mathcal{S}^{(m)})\subset\Omega$ (typically a queue); $\Lambda_{\min}\in\mathbb{R}_{+}$, minimum edge length allowed for sparse grid boxes; $\tau\in(0,1]$, detection tolerance.
@@
4:$C\leftarrow\emptyset$triangleright set of ``visited'' center-edge pairs
@@
14:              if $\lambda\geq\Lambda_{\min}$ and $(\boldsymbol{x}^{\prime}_{i},\lambda)\not\in C$ then
15:                  insert $(\boldsymbol{x}^{\prime}_{i},\lambda)$ in $L$triangleright typically as last element, if $L$ is a queue
16:              else
17:                  $T\leftarrow T\cup\{\boldsymbol{x}_{i}^{\prime}\}$
18:              end if
19:         end if
20:     end for
21:end while
@@
In Algorithm 1 the sparse grid are built on hypercubic boxes. This is not a restricting assumption, as the algorithm can be generalized resorting to sparse grid built on hyperrectangles, changing only the stopping criterion based on the box's edge length. For the sake of simplicity we keep the discussion focused on the case on sparse grids built upon hypercubic boxes.
@@
The algorithm exploits equispaced sparse grids for building a sort of nested structure that, in case of an exact discontinuity detector, can guarantee convergence to the discontinuity interfaces of $g$ (see 3.1, next); moreover, this structure permits to save computational resources because it increases the probability that some points of a new sparse grid, generated at step 3, are shared with other sparse grids. In order to enhance this characteristic it is important that the boxes of the starting sparse grids are centered properly (i.e., according to the collocation knots criterium), if more than one is given at the first step of the algorithm. Theoretically, the algorithm can be extended also to non-equispaced sparse grids, but the generation of the new sparse grids (step 3) cannot be based on the longest edge touching the troubled point, if we want to obtain a behavior similar to the one described above. Then, generalization to non-equispaced sparse grids is postponed to future work.
@@
Algorithm 1 represents the backbone of sparse grid-based algorithms for discontinuity detection. Then, depending on the user's needs, the algorithm can be changed and/or improved, e.g.:
@@
  • the troubled points' detection, for each box in the list $L$, can be optimized and/or parallelized.
@@
It is worth remarking that any NN-based discontinuity detector is particularly suitable to be used with the algorithm for the troubled points predictions (see the last item in 3.4); indeed, NNs can make predictions of thousands of inputs in the order of the second, exploiting the inner operations characterizing the NNs. In order to fully exploit the potential of NNs, we propose a modification of Algorithm 1. In particular, given NN-based detector $\Delta$, we can compute the vectors $\boldsymbol{p}^{(1)},\ldots,\boldsymbol{p}^{(K)}\in\mathbb{R}^{N}$ corresponding to the points of the sparse grids $\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(K)}$, respectively, with just one NN batch-evaluation with respect to the vectors $\boldsymbol{g}^{(1)},\ldots,\boldsymbol{g}^{(K)}$; i.e.,
@@
Therefore, the efficiency of the algorithm increases because we can evaluate entire batches of sparse grids at once and not one-by-one (see lines 4-8 in Algorithm 1). This new NN-based method is detailed in Algorithm 2.
@@
1:$g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$, function; $\mathcal{S}$, sparse grid (equispaced, hypercubic box); $\Delta:\mathbb{R}^{N}\rightarrow[0,1]^{N}$, NN-based discontinuity detector (only-eval. dep.), based on $\mathcal{S}$; $L$, set of center-edge length pairs characterizing the boxes $\mathbb{B}(\mathcal{S}^{(1)}),\ldots,\mathbb{B}(\mathcal{S}^{(m)})\subset\Omega$; $\Lambda_{\min}\in\mathbb{R}_{+}$, minimum edge length allowed for sparse grid boxes; $\tau\in(0,1]$, detection tolerance.
@@
4:$C\leftarrow\emptyset$triangleright set of ``visited'' center-edge pairs
@@
9:     $\Gamma\in\mathbb{R}^{K\times N}\leftarrow$ matrix $(g(\boldsymbol{x}_{i}^{(k)}))$triangleright ``batch of inputs'', see (13)
10:     $P\in[0,1]^{K\times N}\leftarrow$ $\Delta(\Gamma)$triangleright ``batch-predictions'' of $\Delta$, see (12) and (13)
@@
14:              if $\lambda\geq\Lambda_{\min}$ and $(\boldsymbol{x}_{i}^{(k)},\lambda)\not\in C$ then
15:                  $L\leftarrow L\cup\{(\boldsymbol{x}_{i}^{(k)},\lambda)\}$
16:              else
17:                  $T\leftarrow T\cup\{\boldsymbol{x}_{i}^{(k)}\}$
18:              end if
19:         end if
20:     end for
21:end while
We conclude observing the convergence properties of Algorithm 1 under particular assumptions.
Theorem 3.1 (Algorithm Convergence).
Let $\mathcal{S}$ be an equispaced sparse grid of $N>1$ points in $\mathbb{R}^{n}$ with minimum discretization level greater than one, for each dimension. Let $\Delta^{*}$ be an exact discontinuity detector based on $\mathcal{S}$. Let $\delta$ be the set of all and only the discontinuity points of a discontinuous function $g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$ and let $\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(m)}$ be the $m$ starting sparse grids of Algorithm 1.
@@
Without loss of generality, we assume to start the algorithm with a unique sparse grid $\mathcal{S}$. Then, let $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$ be an edge of the SGG based on $\mathcal{S}$ that intersects $\delta$ in $\boldsymbol{x}_{i}^{\delta}$ (see 3.1). Let $\lambda_{i}$ be the length of the longest edge with $\boldsymbol{x}_{i}$ as vertex and let $\Lambda$ be the edge length of the sparse grid box $\mathbb{B}(\mathcal{S})$; then, it holds
$\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\delta}}\|\leq\frac{\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{j}}\|}{2}\leq\frac{\lambda_{i}}{2}<\frac{\Lambda}{2}\,.$ (15)
In particular, $\lambda_{i}$ is strictly smaller than $\Lambda$ due to the hypothesis on the minimum discretization level chosen for $\mathcal{S}$. Moreover, since we are considering only equispaced collocation knots, we have that $\lambda_{i}\leq\Lambda/2$; then, we can rewrite (15) as
$\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\delta}}\|\leq\frac{\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{j}}\|}{2}\leq\frac{\lambda_{i}}{2}\leq\frac{\Lambda}{4}\,.$ (16)
@@
According to Algorithm 1, there are two possibilities: $i$) $\lambda_{i}<\Lambda_{\min}$, then stop and assign $\boldsymbol{x}_{i}$ to $T$; $ii$) $\lambda_{i}\geq\Lambda_{\min}$, then refine. In the first case, (14) holds, because $\mathrm{dist}(\boldsymbol{x}_{i},\delta)=\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\delta}}\|$. In the second case, we consider the sparse grid $\mathcal{S}^{\prime}$, centered on $\boldsymbol{x}_{i}$ and with box's edge length $\Lambda^{\prime}=\lambda_{i}$. Now, let $e^{\prime}_{i^{\prime}j^{\prime}}=\{\boldsymbol{x}^{\prime}_{i^{\prime}},\boldsymbol{x}^{\prime}_{j^{\prime}}\}$ be an edge of the SGG based on $\mathcal{S}^{\prime}$ that intersects $\delta$ in $\boldsymbol{x}^{\prime\delta}_{i^{\prime}}$. Let $\lambda^{\prime}_{i^{\prime}}$ be the length of the longest edge with $\boldsymbol{x}^{\prime}_{i^{\prime}}$ as vertex; then, it holds
$\|{\boldsymbol{x}^{\prime}_{i^{\prime}}-\boldsymbol{x}^{\prime\delta}_{i^{\prime}}}\|\leq\frac{\lambda^{\prime}_{i^{\prime}}}{2}\leq\frac{\Lambda^{\prime}}{4}\leq\frac{\Lambda}{8}\,.$ (17)
We can easily generalize (17) to the $k$-th level of refinement, $k\in\mathbb{N}$; i.e.:
$\|{\boldsymbol{x}^{(k)}-\boldsymbol{x}^{(k)\,\delta}}\|\leq\frac{\lambda^{(k)}}{2}\leq\frac{\Lambda^{(k)}}{4}\leq\frac{\Lambda}{2^{k+2}}\quad\text{and}\quad\lambda^{(k)}\leq\frac{\Lambda}{2^{k+1}}\,,$ (18)
where the index $i^{(k)}$ is omitted for the ease of notation.

### v4

@@ 3 Discontinuity Detection and Sparse Grids
The discontinuity detection method proposed herein aims to extend the one proposed in [3] moving from uniform grids to sparse grids. In view of this extension, we seek for a NN architecture leveraging sparse grids; thus, we consider the use of Graph-Instructed Neural Networks (GINNs) built with respect to a SGG. We remark that the method proposed herein is designed in such a way that also classic NNs, such as Multi-Layer Perceptrons (MLPs), can be used; however, the numerical experiments show that the method benefits from the use of GINN (see Section 5); this is in line with the results presented in [26].
For a given function $g:\Omega\subseteq\mathbb{R}^{n}\to\mathbb{R}$, the method proposed relies on the following ingredients:
  • the concept of trouble points of a sparse grid (i.e, points of the grid that are near to a discontinuity interface of $g$);
@@
The algorithm defined in this work does not depend on the nature of the discontinuity detector used, and the function $g$ can be rather general; however, the algorithm works better with discontinuous functions characterized by discontinuity interfaces with small codimensions (with respect to $\mathbb{R}^{n}\supseteq\Omega$). In the following, we formally introduce the notions of troubled point and discontinuity detector based on a sparse grid. The definition of troubled points is inspired by the definition of troubled cells given in [3].
@@
Let $\mathcal{S}$ be a sparse grid in $\mathbb{R}^{n}$ and let $G=(\mathcal{S},E)$ be the (possibly weighted) graph based on $\mathcal{S}$. Let $g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a given function such that $\mathbb{B}(\mathcal{S})\subseteq\Omega$ and let $\delta$ be the set of all and only the discontinuity points of $g$. Then, the node $\boldsymbol{x}_{i}$ of $\mathcal{S}$ is said to be a troubled point with respect to $g$ if there exist $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}\in E$ such that:
  • $\delta\cap\overline{(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}\neq\emptyset\,$;
@@
Let $\mathcal{S}$ be a sparse grid in $\mathbb{R}^{n}$ made of $N$ points and let $\mathcal{S}_{/\simeq}$ denotes the set of all the sparse grids similar to $\mathcal{S}$. Let $\mathcal{F}$ denotes the set of all the functions $g:\Omega\rightarrow\mathbb{R}$, for an arbitrary $\Omega\subseteq\mathbb{R}^{n}$. Then, a discontinuity detector based on $\mathcal{S}$ is a function $\Delta$ that, for each $\mathcal{S}^{\prime}\simeq\mathcal{S}$ and each function $g\in\mathcal{F}$, returns a vector $\boldsymbol{p}=(p_{1},\ldots,p_{N})$, with $0\leq p_{i}\leq 1$ for all $i=1,\ldots,N$, where $p_{i}$ gives an indication of how likely $\boldsymbol{x}_{i}^{\prime}\in\mathcal{S}^{\prime}$ is a troubled point. The closer $p_{i}$ is to $1$, the more likely $\boldsymbol{x}^{\prime}_{i}$ is a troubled point.
According to Definition 3.2, a discontinuity detector provides, for each grid point, a sort of likelihood for being a troubled point. A point $\boldsymbol{x}_{i}^{\prime}$ is labeled as troubled if $p_{i}\geq\tau$, where $\tau\in(0,1]$ is a given threshold.
The previous definition is intentionally qualitative, as several types of discontinuity detectors can be defined. For example, we can define detectors that evaluate $\boldsymbol{p}$ knowing: $i$) the function $g$ and the coordinates of the points in $\mathcal{S}^{\prime}$; $ii$) the values of $g$ taken at the points in $\mathcal{S}^{\prime}$ and the points’ coordinates; $iii$) the values of $g$ taken at the points in $\mathcal{S}^{\prime}$, only. In this work, we focus on building NN-based detectors of the latter type; i.e., detectors $\Delta$ such that $\Delta:(\mathbb{R}\cup\{\infty\})^{N}\rightarrow[0,1]^{N}$ and
@@
However, an exact detector is typically defined through expensive algorithms and in general it is not a detector based on the discontinuous function evaluations only (as in (9)). Therefore, given a sparse grid, our idea consists in building a non-expensive but effective discontinuity detector that is only-evaluations dependent, training a NN model for approximating the exact one; given such a kind of NN-based detector, we can use it with a sparse grid-based discontinuity detection algorithm, aiming to reach very good results at a consistently lower computational cost, even in domains of dimension $n>3$.
Figure 3: Example of sparse grid points classification made for a function $g:\mathbb{R}^{2}\to\mathbb{R}$ with an exact discontinuity detector $\Delta^{*}$ (left) and a inexact discontinuity detector $\Delta$ (right). The black dotted line is the function’s discontinuity interface. Green circles are true troubled points, magenta circles are false troubled points, magenta dots are false non-troubled points. For the inexact detector, we report also the values $p_{i}$ corresponding to the highlighted sparse grid points (threshold $\tau=0.5$).
@@
Similarly to what done in [3], the introduction of a NN-based detector that is only-evaluations dependent (see (9)) is useful to make the detector universally applicable to any sparse grid $\mathcal{S}^{\prime}\simeq\mathcal{S}$ and any discontinuous function with domain dimension equal to $n$. On the other hand, since these detector will not be exact, we can expect that the smaller are the ratios between the function evaluations and the edge lengths (i.e., the slopes $|g(\boldsymbol{x}_{i}^{\prime})-g(\boldsymbol{x}_{j}^{\prime})|/\|{\boldsymbol{x}_{i}^{\prime}-\boldsymbol{x}_{j}^{\prime}}\|$), the lower is the reliability of the predicted troubled points.
@@ 3.1 Algorithm for Discontinuity Detection
We adopt a similar approach for our sparse grid-based method. However, thanks to Proposition 2.1 and Remark 3.1, we can use the same detector $\Delta$ for both the coarse step and any refinement step. Therefore, we define an iterative approach that stops when a minimum box size for the sparse grid is reached. We point the reader to the fact that we can apply several refinements leveraging the use of sparse grids, which allow to save on the number of function evaluations, thus delaying the curse of dimensionality phenomenon.
@@
    using $\Delta$, detect the troubled points of one sparse grid and mark the grid as “checked”;
@@
    for each troubled point $\boldsymbol{x}$, generate a new sparse grid $\mathcal{S}^{\prime}$ similar to $\mathcal{S}$, with box $\mathbb{B}(\mathcal{S}^{\prime})$ centered in $\boldsymbol{x}$ and box’s edge length equal to the length of the longest graph’s edge intersecting the troubled point. If this latter length is shorter than a chosen value $\Lambda_{\min}\in\mathbb{R}_{+}$, do not create the new sparse grid.
@@
    repeat steps 2 and 3, until all the sparse grids are “checked”.
@@
1:$g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$, function; $\mathcal{S}$, sparse grid (equispaced, hypercubic box); $\Delta$, discontinuity detector based on $\mathcal{S}$; $L=((\boldsymbol{x}_{1},\lambda_{1}),\ldots,(\boldsymbol{x}_{m},\lambda_{m}))$, sequence of center and edge length pairs characterizing the sparse grid boxes $\mathbb{B}(\mathcal{S}^{(1)}),\ldots,\mathbb{B}(\mathcal{S}^{(m)})\subset\Omega$ (typically a queue); $\Lambda_{\min}\in\mathbb{R}_{+}$, minimum edge length allowed for sparse grid boxes; $\tau\in(0,1]$, detection tolerance.
@@
4:$C\leftarrow\emptyset$triangleright set of “visited” center-edge pairs
@@
14:              if $\lambda\geq\Lambda_{\min}$ then
15:                  if $(\boldsymbol{x}^{\prime}_{i},\lambda)\not\in C$ then
16:                       insert $(\boldsymbol{x}^{\prime}_{i},\lambda)$ in $L$triangleright typically as last element, if $L$ is a queue
17:                  end if
18:              else
19:                  $T\leftarrow T\cup\{\boldsymbol{x}_{i}^{\prime}\}$
20:              end if
21:         end if
22:     end for
23:end while
@@
In Algorithm 1 the sparse grid are built on hypercubic boxes. This is not a restricting assumption, as the algorithm can be generalized resorting to sparse grid built on hyperrectangles, changing only the stopping criterion based on the box’s edge length. For the sake of simplicity we keep the discussion focused on the case on sparse grids built upon hypercubic boxes.
@@
The algorithm exploits equispaced sparse grids for building a sort of nested structure that, in case of an exact discontinuity detector, can guarantee convergence to the discontinuity interfaces of $g$ (see Theorem 3.1, next); moreover, this structure permits to save computational resources because it increases the probability that some points of a new sparse grid, generated at step 3, are shared with other sparse grids. In order to enhance this characteristic it is important that the boxes of the starting sparse grids are centered properly (i.e., according to the collocation knots criterium), if more than one is given at the first step of the algorithm. Theoretically, the algorithm can be extended also to non-equispaced sparse grids, but the generation of the new sparse grids (step 3) is not straightforward and is postponed to future work.
@@
Algorithm 1 represents the backbone of sparse grid-based algorithms for discontinuity detection. Then, depending on the user’s needs, the algorithm can be changed and/or improved, e.g.:
@@
  • the troubled points’ detection, for each box in the list $L$, can be optimized and/or parallelized.
@@
It is worth remarking that any NN-based discontinuity detector is particularly suitable to be used with the algorithm for the troubled points predictions (see the last item in Remark 3.4); indeed, NNs can make predictions of thousands of inputs in the order of the second, exploiting the inner operations characterizing the NNs. In order to fully exploit the potential of NNs, we propose a modification of Algorithm 1. In particular, given NN-based detector $\Delta$, we can compute the vectors $\boldsymbol{p}^{(1)},\ldots,\boldsymbol{p}^{(K)}\in\mathbb{R}^{N}$ corresponding to the points of the sparse grids $\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(K)}$, respectively, with just one NN batch-evaluation with respect to the vectors $\boldsymbol{g}^{(1)},\ldots,\boldsymbol{g}^{(K)}$; i.e.,
@@
Therefore, the efficiency of the algorithm increases because we can evaluate entire batches of sparse grids at once and not one at a time (see lines 4-8 in Algorithm 1). This new NN-based method is detailed in Algorithm 2.
@@
1:$g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$, function; $\mathcal{S}$, sparse grid (equispaced, hypercubic box); $\Delta:\mathbb{R}^{N}\rightarrow[0,1]^{N}$, NN-based discontinuity detector (only-eval. dep.), based on $\mathcal{S}$; $L$, set of center and edge length pairs characterizing the boxes $\mathbb{B}(\mathcal{S}^{(1)}),\ldots,\mathbb{B}(\mathcal{S}^{(m)})\subset\Omega$; $\Lambda_{\min}\in\mathbb{R}_{+}$, minimum edge length allowed for sparse grid boxes; $\tau\in(0,1]$, detection tolerance.
@@
4:$C\leftarrow\emptyset$triangleright set of “visited” center-edge pairs
@@
9:     $\Gamma\in\mathbb{R}^{K\times N}\leftarrow$ matrix $(g(\boldsymbol{x}_{i}^{(k)}))$triangleright “batch of inputs”, see (11)
10:     $P\in[0,1]^{K\times N}\leftarrow$ $\Delta(\Gamma)$triangleright “batch-predictions” of $\Delta$, see (10) and (11)
@@
14:              if $\lambda\geq\Lambda_{\min}$ then
15:                  if $(\boldsymbol{x}_{i}^{(k)},\lambda)\not\in C$ then
16:                       $L\leftarrow L\cup\{(\boldsymbol{x}_{i}^{(k)},\lambda)\}$
17:                  end if
18:              else
19:                  $T\leftarrow T\cup\{\boldsymbol{x}_{i}^{(k)}\}$
20:              end if
21:         end if
22:     end for
23:end while
We conclude by observing the convergence properties of Algorithm 1 under particular assumptions.
Theorem 3.1 (Algorithm Convergence with Exact Detector).
Let $\mathcal{S}$ be an equispaced sparse grid of $N>1$ points in $\mathbb{R}^{n}$ with minimum refinement level greater than one, for each dimension. Let $\Delta^{*}$ be an exact discontinuity detector based on $\mathcal{S}$. Let $\delta$ be the set of all and only the discontinuity points of a discontinuous function $g:\Omega\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$ and let $\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(m)}$ be the $m$ starting sparse grids of Algorithm 1.
@@
Without loss of generality, we assume to start the algorithm with a unique sparse grid $\mathcal{S}$. Let $\Lambda$ denote the edge length of the sparse grid box $\mathbb{B}(\mathcal{S})$. Let $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$ be an edge of the SGG based on $\mathcal{S}$ that intersects $\delta$ and, without loss of generality, let us assume that $\boldsymbol{x}_{i}$ is a troubled point (see Definition 3.1); we denote by $\boldsymbol{x}_{\delta}$ the point in $\delta\cap e_{ij}$ nearest to $\boldsymbol{x}_{i}$. Let $\lambda_{i}$ be the length of the longest edge with $\boldsymbol{x}_{i}$ as vertex; then, it holds
$\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{\delta}}\|\leq\frac{\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{j}}\|}{2}\leq\frac{\lambda_{i}}{2}.$ (13)
Since we assume that at least one refinement is made for each dimension, and since we use equispaced nodes as initial univariate distribution, we have $\lambda_{i}\leq\Lambda/2$; hence we can rewrite (13) as
$\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{\delta}}\|\leq\frac{\Lambda}{4}\,.$ (14)
According to Algorithm 1, two scenarios are possible: ($i$) $\lambda_{i}<\Lambda_{\min}$, then stop and assign $\boldsymbol{x}_{i}$ to $T$; ($ii$) $\lambda_{i}\geq\Lambda_{\min}$, then refine.
If case ($i$) occurs, (12) holds true because
$\mathrm{dist}(\boldsymbol{x}_{i},\delta)\leq\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{\delta}}\|\leq\lambda_{i}/2<\Lambda_{\min}/2\,\leq\,\varepsilon.$
@@
If case ($ii$) occurs, consider the sparse grid $\mathcal{S}^{\prime}$ introduced in the refinement step, centered on $\boldsymbol{x}_{i}$ and with box edge length $\Lambda^{\prime}=\lambda_{i}$. Let $e^{\prime}_{pq}=\{\boldsymbol{x}^{\prime}_{p},\boldsymbol{x}^{\prime}_{q}\}$ be an edge of the SGG based on $\mathcal{S}^{\prime}$ that intersects $\delta$ such that, without loss of generality, $\boldsymbol{x}^{\prime}_{p}$ is a troubled point; we denote by $\boldsymbol{x}^{\prime}_{\delta}$ the point in $\delta\cap e^{\prime}_{pq}$ nearest to $\boldsymbol{x}^{\prime}_{p}$. Let $\lambda^{\prime}_{p}$ be the length of the longest edge with $\boldsymbol{x}^{\prime}_{p}$ as vertex; then, it holds
$\|{\boldsymbol{x}^{\prime}_{p}-\boldsymbol{x}^{\prime}_{\delta}}\|\leq\frac{\lambda^{\prime}_{p}}{2}\leq\frac{\Lambda^{\prime}}{4}\leq\frac{\Lambda}{8}$ (15)
where the third inequality is due to the fact that ${\Lambda^{\prime}}\leq\frac{\Lambda}{2}$ because at least one refinement is made, and the second inequality applies because ${\lambda^{\prime}_{p}}\leq\frac{\Lambda^{\prime}}{2}$, as any edge length is at least half of the sparse grid box.
Repeating the same reasoning, we obtain after $k$ refinement steps, $k\geq 1$, the following inequalities:
$\|{\boldsymbol{x}^{(k)}-\boldsymbol{x}^{(k)}_{\delta}}\|\leq\frac{\lambda^{(k)}}{2}\leq\frac{\Lambda^{(k)}}{4}\leq\frac{\Lambda}{2^{k+2}}\quad\text{and}\quad\lambda^{(k)}\leq\frac{\Lambda}{2^{k+1}}\,,$ (16)
where the subscript for $\boldsymbol{x}^{(k)},\lambda^{(k)}$ is omitted for the ease of notation.